OpenStudy (alibaby):

Which of the following has a graph that is a straight line? Equation 1: y = 2x + 7 Equation 2: y2 = x − 1 Equation 3: y = 2x2 + 4 Equation 4: y = 3x3

OpenStudy (jailbird):

Use your knowledge to eliminate, what happens is you raise the x to a power??

OpenStudy (alibaby):

@Awolflover1 @Hayhayz @SolomonZelman

OpenStudy (alibaby):

idk

OpenStudy (jailbird):

It curves...

OpenStudy (alibaby):

im not sure that's why im on here asking

OpenStudy (jailbird):

its ok, when you raise x to a power, it curves up or down depending on whether its positive or negative, raising y to a power is going to cause it to curve to the left or right, which means???

OpenStudy (jailbird):

Does that help??

OpenStudy (alibaby):

nope

OpenStudy (jailbird):

If a graph curves, is it straight?

OpenStudy (alibaby):

no

OpenStudy (alibaby):

@mathmale

OpenStudy (jailbird):

and x raised to a power makes it curve up and down, if it curves up and down is it straight??

OpenStudy (alibaby):

this is very confusing

OpenStudy (mathstudent55):

An equation has a straight line if its variables are alone or just multiplied by numbers. If it has variables in the denominator, variables raised to a power, variables multiplied by variables, variables in roots, logarithms of variables, or other non-linear functions with variables, then its graph is not a straight line.

OpenStudy (alibaby):

i think it's A or D is it one of those

OpenStudy (mathmale):

First off, " Equation 2: y2 = x − 1" should be written as y^2 = x − 1. That " ^2 " indicates the "squaring function," which you've likely seen before. Likewise, Equations 3 and 4 should be written y = 2x^2 + 4 and y = 3x^3 Here is how we classify these functions that involve powers of x: x^0: constant function; graph is a horizontal line x^1: linear function; graph is a straight line x^2: quadratic function; graph is a parabola opening up x^3: cubing function: graph is odd, begins in the 3rd quadrant, goes thru the origin and ends up in the 1st quadrant. This may be more info that you were looking for, but trust me, you do need to know and be able to recognize these kinds of functions and their graphs.

OpenStudy (mathmale):

With this info, can you now identify the function that has a straight line graph? Equation 1: y = 2x + 7 Equation 2: y^2 = x − 1 Equation 3: y = 2x^2 + 4 Equation 4: y = 3x^3

OpenStudy (alibaby):

im sooo confused -banging head on desk-

OpenStudy (mathmale):

I'm sorry you're feeling confused. But there is basic info in every subject that we (meaning me as well as you) MUST learn.

OpenStudy (mathmale):

Hint: x^1 and x are exactly the same.

OpenStudy (mathmale):

If the highest power of x is 1, then the function is "linear" and the graph is a straight line.

OpenStudy (mathstudent55):

Examples 1. Linear functions whose graphs are straight lines: \(y = 2x + 5\) \(2x + 5y = -2\) \(y = -\sqrt 2 x\) <--- notice only the 2 is in the square root, not the x \(\dfrac{1}{2} x + \dfrac{3}{4}y = \dfrac{\sqrt{55}}{21} \) 2. Non-linear functions whose graphs are not straight lines: \(\dfrac{1}{x} = 5y\) <--- x in denominator \(\sqrt{x + 4} = 2y - 5\) <--- x in root \(2xy + 3x = 5y\) <--- product of variables \(y = 2x^2 - 4\) <--- variable raised to a power \(y = \log (2x) \) <--- logarithm of a variable

OpenStudy (alibaby):

is it B?

OpenStudy (phi):

a line has *no exponents" on the variables which equation has no "little numbers" on the letters

OpenStudy (mathmale):

Please, Ali, explain your reasoning.

OpenStudy (alibaby):

i think it's B because it's a straight line it dosent curve

OpenStudy (mathmale):

Right, and it's a straight line because of what else? Hint: Look at the exponent of x in the given equation.

OpenStudy (phi):

There is no B. Do you mean Equation 2: \(y^2 = x − 1\) ?

OpenStudy (alibaby):

yes

OpenStudy (alibaby):

but

OpenStudy (phi):

a line has *no exponents" on the variables an example of an exponent is the 2 next to the y

OpenStudy (alibaby):

B's right?

OpenStudy (mathmale):

What I see in Answer B is y^2=x-1. Did you check the exponent of x in this equation? If not, please do so now.

OpenStudy (phi):

do you see the \( y^2 \) ? that makes it *NOT* a line

OpenStudy (mathmale):

I should have said that if you see a power other than 1 on either x or y, the function is NOT linear.

OpenStudy (mathmale):

Once again, pls review all four possible answers. Only one of them features an exponent of 1 on both x and y.

OpenStudy (mathmale):

...or, no exponent at all. Earlier I explained to you that 1x^1 is the same as 1x or just x.